OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

hartnn (hartnn):

do you know how to subtract 2 equations ? if you know, subtract those 2 equations, and 'y' will get eliminated! :)

OpenStudy (anonymous):

won't i have to do the opposite

hartnn (hartnn):

what does that mean ? if that means making the sign of 'y' as opposite, so that you can eliminate it, then yes :) subtracting will make the sign of y opposite.

OpenStudy (anonymous):

Can you walk me through

hartnn (hartnn):

sure, subtracting 2 equations means subtracting left sides and right sides of them \(\Large (-2x+ y ) - (-4x+y) = 1 - (-1)\) can you try to simplify that ^ ?

OpenStudy (anonymous):

well what i would do is take one of the equations and turing it into y=mx+b form, and then substitute it into the other equation. so you could change -4x+y=-1 and write it as y=4x-1 then substitute into the other equation so your doing -2x+(4x-1)=1 then just simplify.if you do this correctly you should be left with 2x-1=1. get x by itself, giving you x=1. plug x back into y=4x-1 and you are left with y=3. your answer would be Y=3 and X=1

OpenStudy (anonymous):

thats using substitution method

OpenStudy (anonymous):

and @hartnn i dont get you method

hartnn (hartnn):

thats called elimination method. if you observe, 'y' gets eliminated and you directly get the value of x

OpenStudy (anonymous):

your way was easiest @hartnn It took me awhile but i got it using both.

hartnn (hartnn):

i am glad you got it :)

OpenStudy (anonymous):

but for equations like that dont you have to find the value for both x and y?

hartnn (hartnn):

ofcourse! once you get 'y', use any of the 2 original equations to get the value of x :)

hartnn (hartnn):

oh, here you got x then use any of the 2 original equations to get the value of y :)

OpenStudy (anonymous):

i see that makes sense. thank you that is easier it seems. again thank you.

hartnn (hartnn):

and welcome ^_^